# Algorithmic Decision-Making under Agents with Persistent Improvement

### v1

@@ Abstract
This paper studies algorithmic decision-making under human’s strategic behavior, where a decision maker uses an algorithm to make decisions about human agents, and the latter with information about the algorithm may exert effort strategically and improve to receive favorable decisions. Unlike prior works that assume agents benefit from their efforts immediately, we consider realistic scenarios where the impacts of these efforts are persistent and agents benefit from efforts by making improvements gradually. We first develop a dynamic model to characterize persistent improvements and based on this construct a Stackelberg game to model the interplay between agents and the decision-maker. We analytically characterize the equilibrium strategies and identify conditions under which agents have incentives to improve. With the dynamics, we then study how the decision-maker can design an optimal policy to incentivize the largest improvements inside the agent population. We also extend the model to settings where 1) agents may be dishonest and game the algorithm into making favorable but erroneous decisions; 2) honest efforts are forgettable and not sufficient to guarantee persistent improvements. With the extended models, we further examine conditions under which agents prefer honest efforts over dishonest behavior and the impacts of forgettable efforts.
@@ 1 Introduction
In applications such as lending, college admission, hiring, etc., machine learning (ML) algorithms have been increasingly used to evaluate and make decisions about human agents. Given information about an algorithm, agents subject to ML decisions may behave strategically to receive favorable decisions. How to characterize the strategic interplay between algorithmic decisions and agents, and analyze the impacts they each have on the other, are of great importance but challenging.
@@
This paper studies algorithmic decision-making under strategic agent behavior. Specifically, we consider a decision-maker who assesses a group of agents and aims to accept those that are qualified for certain tasks based on assessment outcomes. With knowledge of the acceptance rule, agents may behave strategically to increase their chances of getting accepted. For example, agents may invest to genuinely improve their qualifications (i.e., honest effort), or they may manipulate the observable assessment outcomes to game the algorithm (i.e., dishonest effort). Both types of behaviors have been studied. In particular, Hardt et al. (2016a); Dong et al. (2018); Braverman and Garg (2020); Jagadeesan et al. (2021); Sundaram et al. (2021); Zhang et al. (2022); Eilat et al. (2022) focus on learning under strategic manipulation, where they proposed various analytical frameworks (e.g., Stackelberg games) to model manipulative behavior, and analyzed models or developed learning algorithms that are robust against manipulation.
@@
Another line of research (Zhang et al., 2020; Harris et al., 2021; Bechavod et al., 2022; Kleinberg and Raghavan, 2020; Chen et al., 2020; Barsotti et al., 2022; Jin et al., 2022) considers a different setting where agent qualifications (labels) change in accordance with the improvement actions. The goal of the decision-maker is to design a mechanism such that agents are incentivized to behave toward directions that improve the underlying qualifications. Notably, Kleinberg and Raghavan (2020) proposed a mechanism to incentivize individuals to invest in specific improvable features. Their work inherited the classical settings of the Principal-agent model in economics but designed an incentivizing mechanism under a linear classifier. They modeled manipulation and improvement similarly (linear in efforts) and did not consider the persistent and delayed effects of improvement. The mixture of both improvement and manipulation behavior is also studied (Miller et al., 2020; Chen et al., 2020; Barsotti et al., 2022; Horowitz and Rosenfeld, 2023). However, these works regarded improvement as a similar action to manipulation where the only difference is it will incur a label change. Another related topic is performative prediction (Perdomo et al., 2020; Izzo et al., 2021; Hardt et al., 2022; Jin et al., 2024), an abstraction that captures agent actions via model-induced distribution shifts. Details and more related works are presented in Appendix C.
@@
This paper primarily focuses on honest agents with improvement, while settings with both improvement and manipulation are also studied. We first propose a novel two-stage Stackelberg game to model the interactions between decision-maker and agents, i.e., the decision-maker commits to its policy, following which agents best respond. A crucial difference between this study and the prior works is that the existing models all assume that the results of agents’ improvement actions are immediate, i.e., once agents decide to improve, they experience sudden changes in qualifications and receive the return at once. However, we observe that in many real-world applications, the impacts of improvement action are indeed persistent and delayed. For example, humans improve their abilities by acquiring new knowledge, but they make progress gradually and benefit from such behavior throughout their lifetime; loan applicants improve their credit behaviors by repaying all the debt in time, but there is a time lag between such behaviors and the increase in their credit scores. Therefore, it is critical to capture these delayed outcomes in the Stackelberg game formulation.
@@ 2 Problem Formulation
Decision-maker’s policy. For an agent with qualification profile $q_{t}$, the decision-maker assesses whether the agent’s profile lines up with the desired qualifications $d$, and makes decision $D_{t}$ based on their similarity $x_{t}:=q_{t}^{T}d$ using a fixed threshold policy $\pi(x_{t})=\textbf{1}(x_{t}\geq\theta)$, i.e., only agents that are sufficiently fit can get accepted. How to choose threshold $\theta$ is discussed in Sec. 4. We assume only agents with initial similarity $x_{0}\geq 0$ are interested in positions and only focus on these candidates.
Although the decision policy introduced above focuses on the similarity between $q_{t}$ and $d$ where qualifications $q_{t}$ are normalized with the same magnitude for all agents, it can be easily extended to settings where the magnitude/strength of skills also matters and may differ across agents. Specifically, we propose a pre-normalization procedure to account for the strength of skills. The idea is to first add an additional dimension to initial qualification profile $q_{0}$, which represents the agent’s unobservable “irrelevant attribute" (all other skills an agent has that are not important for the decision). Meanwhile, we add this dimension to ideal qualification profile $d$ with 0 as its value. We can make a natural assumption that after adding the dimension of “irrelevant attribute", the norm of $(m+1)$-dimensional complete profile is the same for all agents. This is reasonable and supported by the literature (Liu et al., 2022; Holmstrom and Milgrom, 1991), which suggests that when qualification profiles are multi-dimensional, the competency in relevant/measurable attributes implies the weakness in irrelevant/unmeasurable attributes. The detailed pre-normalization procedure is formally presented in Algorithm 1 (App. A).
Agent qualification dynamics. We assume agents have information about the ideal profile $d$ (e.g., from application guides, mock interviews). In the beginning, agents with $q_{0}$ can choose to improve their profiles by investing an effort $k\in[0,1]$ to acquire the relevant knowledge, but the effort will have delayed and persistent effects over subsequent time stages. The specific value of $k$ depends on the agent’s utility and will be introduced at the end of this section. Upon investing $k$, the agent’s qualifications $q_{t}$ gradually improve over time based on the following:
@@
Dynamics in (1) model the delayed and persistent impacts of improvement action (i.e., effort $k$). In many real applications, humans acquire knowledge and benefit from repeated practices. They make progress toward the goal gradually, and it takes time to receive the desired outcome from the investment. Indeed, (1) is inspired by the dynamics in Dean and Morgenstern (2022), which was used for modeling preference shifts (details are in App. C.3), where individuals update their opinions/preferences based on their correlations with some influencer (e.g., a political figure) and control the power of the intervention. We believe this is similar to improvement especially when agents improve themselves by imitating some "role models". For example, in a job application scenario, the “influencer” is a current worker who holds information session and introduces her profile $d$. Then the agents strive to mimic the profile of $d$ by updating $q_{t}$. The imitating nature of improvement is well justified in many works (e.g., Raab and Liu (2021); Zhang et al. (2022)), making agent improvement suitable to be modeled in a similar fashion to concept drift/preference shift. Thus, we use (1) to model the evolution of agents’ (pre-normalized) qualifications. Based on Prop. 1 of Dean and Morgenstern (2022), we know that $q_{t}$ converges under dynamics, as formally stated in Lemma 2.1 below.
@@
Lemma 2.1 suggests that any agent eventually becomes an ideal candidate with a perfectly aligned profile (i.e., $x_{t}=q_{t}^{T}d=1$), as long as he/she is interested in the position ($x_{0}\geq 0$) and willing to make an effort ($k>0$). The only difference among agents is the speed of convergence: it takes less time for agents who are more qualified at the beginning (i.e., larger $x_{0}$) and/or make more effort (i.e., larger $k$) to become ideal and get accepted. Note that our work focuses on agent’s improvement behavior with persistent and delayed effects. Although the model is presented in a simplified setting where only a one-step effort is made by the agents at the beginning, it can capture more complicated scenarios where agents repeatedly exert efforts multiple times until they reach the target. Each effort has persistent effects on improving the qualification. Specifically, suppose each agent at time $t$ can exert an effort $k_{t}$ and the agent’s qualification improves based on $\widetilde{q}_{t+1}=q_{t}+\sum_{\tau=0}^{t}k_{\tau}\cdot q_{t}^{T}d\cdot d$ with $q_{t+1}=\frac{\widetilde{q}_{t+1}}{\|\widetilde{q}_{t+1}\|_{2}}$ (i.e., every time the agent improves from all accumulated efforts $\sum_{\tau=0}^{t}k_{\tau}\in[0,1]$ he/she invested so far). For this new dynamics, the overall impacts of these efforts $\{k_{t}\}_{t\geq 0}$ on improving agent qualification can be equivalently characterized by the dynamics (1) with some one-step effort. That is, there exists an effort $k^{*}\in[0,1]$ such that investing $k^{*}$ once at the beginning has the same impact on $\lim_{t\to\infty}q_{t}$ as investing a sequence of efforts $\{k_{t}\}_{t\geq 0}$ over time. We provide more detailed discussion on this in App. B. We also discuss a special case where the effect of $k$ is decreasing over time and provide further convergence analysis in App. B.
Agent’s utility & action. Because it takes time for agents to receive rewards (i.e., get accepted) for their efforts, they may not have incentives to invest if there is a long delay. In practice, people may be more attracted to investments with immediate rewards than delayed rewards, or they may simply not have enough time to wait. For example, students only have limited time to prepare for college applications; credit card applicants may not have incentives to improve their credit scores and wait to get approval for a specific credit card when there are many instant-approval cards on the market.
@@
That is, the utility is the exponentially discounted reward an agent receives from the acceptance minus the effort. $r>0$ is the discounting factor. Note that the discounted utility model [^1] Under exponential discounting function, the agent’s reward diminishes at a constant rate (Grüne-Yanoff, 2015). Our model can also adopt other discounting functions (e.g., hyperbolic discounting) for settings when the agent’s reward decreases inconsistently. The qualitative results of this paper still remain the same. has been widely used in literature such as reinforcement learning (Kaelbling et al., 1996), finance (Meier and Sprenger, 2013), and economics (Krahn and Gafni, 1993; Samuelson, 1937).
@@
Stackelberg game. We model the strategic interplay between the decision-maker and agents as a Stackelberg game, which consists of two stages: (i) the decision-maker first publishes the optimal acceptance threshold $\theta$ (details are in Sec. 4); (ii) agents after observing the threshold take actions to maximize their utilities as given in (5).
Manipulation & forgetting. The model formulated above has two implicit assumptions: (i) agents are honest and they improve their qualifications by making actual efforts; (ii) once agents make a one-time effort $k$ to acquire the knowledge, they never forget and can repeatedly leverage this knowledge to improve their profiles based on (1). However, these assumptions may not hold. In practice, agents may fool the decision-maker by directly manipulating $x_{t}$ to get accepted without improving actual $q_{t}$, e.g., people cheat on exams or interviews to get accepted. Moreover, the knowledge agents acquired at the beginning may not be sufficient to ensure repeated improvements.
To capture these, we further extend the above model to two settings:
@@
Objective. In this paper, we study the above interactions between decision-maker and agents. We aim to understand (i) under what conditions agents have incentives to improve their qualifications; (ii) how to design the optimal policy to incentivize the largest improvements inside the agent population; (iii) how the agents would behave when they have both options of manipulation and improvement, and under what conditions agents prefer improvement over manipulation; (iv) how the forgetting mechanism affects agent’s behavior and long-term qualifications.

### v2

@@ Abstract
This paper studies algorithmic decision-making under human strategic behavior, where a decision maker uses an algorithm to make decisions about human agents, and the latter with information about the algorithm may exert effort strategically and improve to receive favorable decisions. Unlike prior works that assume agents benefit from their efforts immediately, we consider realistic scenarios where the impacts of these efforts are persistent and agents benefit from efforts by making improvements gradually. We first develop a dynamic model to characterize persistent improvements and based on this construct a Stackelberg game to model the interplay between agents and the decision-maker. We analytically characterize the equilibrium strategies and identify conditions under which agents have incentives to invest efforts to improve their qualifications. With the dynamics, we then study how the decision-maker can design an optimal policy to incentivize the largest improvements inside the agent population. We also extend the model to settings where 1) agents may be dishonest and game the algorithm into making favorable but erroneous decisions; 2) honest efforts are forgettable and not sufficient to guarantee persistent improvements. With the extended models, we further examine conditions under which agents prefer honest efforts over dishonest behavior and the impacts of forgettable efforts.
@@ 1 Introduction
In applications such as lending, college admission, hiring, recommendation systems, etc., machine learning (ML) algorithms have been increasingly used to evaluate and make decisions about human agents. Given information about an algorithm, agents subject to ML decisions may behave strategically to receive favorable decisions. How to characterize the strategic interplay between algorithmic decisions and agents, and analyze the impacts they each have on the other, are of great importance but challenging
@@
This paper studies algorithmic decision-making under strategic agent behavior. Specifically, we consider a decision-maker who assesses a group of agents and aims to accept those that are qualified for certain tasks based on assessment outcomes. With knowledge of the acceptance rule, agents may behave strategically to increase their chances of getting accepted. For example, agents may invest to genuinely improve their qualifications (i.e., honest effort), or they may manipulate the observable assessment outcomes to game the algorithm (i.e., dishonest effort). Both types of behaviors have been studied. In particular, Hardt et al. (2016a); Dong et al. (2018); Braverman and Garg (2020); Jagadeesan, Mendler-Dünner, and Hardt (2021); Sundaram et al. (2021); Zhang et al. (2022); Eilat et al. (2022) focus on learning under strategic manipulation, where they proposed various analytical frameworks (e.g., Stackelberg games) to model manipulative behavior, and analyzed models or developed learning algorithms that are robust against manipulation.
@@
Another line of research (Zhang et al. 2020; Harris, Heidari, and Wu 2021; Bechavod et al. 2022; Kleinberg and Raghavan 2020; Chen, Wang, and Liu 2020; Barsotti, Kocer, and Santos 2022; Jin et al. 2022) considers a different setting where agent qualifications (labels) change in accordance with the improvement actions. The goal of the decision-maker is to design a mechanism such that agents are incentivized to behave toward directions that improve the underlying qualifications. Notably, Kleinberg and Raghavan (2020) proposed a mechanism to incentivize individuals to invest in specific improvable features. Their work inherited the classical settings of the Principal-agent model in economics but designed an incentivizing mechanism under a linear classifier. They modeled manipulation and improvement similarly (linear in efforts) and did not consider the persistent and delayed effects of improvement. The mixture of both improvement and manipulation behavior is also studied (Miller, Milli, and Hardt 2020; Chen, Wang, and Liu 2020; Barsotti, Kocer, and Santos 2022; Horowitz and Rosenfeld 2023; Jin et al. 2022). However, these works regarded improvement as a similar action to manipulation where the only difference is it will incur a label change. Another related topic is performative prediction (Perdomo et al. 2020; Izzo, Ying, and Zou 2021; Hardt, Jagadeesan, and Mendler-Dünner 2022; Jin et al. 2024b), an abstraction that captures agent actions via model-induced distribution shifts. Details and more related works are presented in Sec. 2.
@@
This paper primarily focuses on honest agents (i.e., agents will invest effort to improve their qualifications), while settings with both improvement and manipulation are also studied. We first propose a novel two-stage Stackelberg game to model the interactions between decision-maker and agents, i.e., the decision-maker commits to its policy, following which agents best respond. A crucial difference between this study and the prior works is that the existing models all assume that the results of agents’ improvement actions are immediate, i.e., once agents decide to improve, they experience sudden changes in qualifications and receive the return at once. However, we observe that in many real-world applications, the impacts of improvement action are indeed persistent and delayed. For example, humans improve their abilities by acquiring new knowledge, but they make progress gradually and benefit from such behavior throughout their lifetime; loan applicants improve their credit behaviors by repaying all the debt in time, but there is a time lag between such behaviors and the increase in their credit scores. Therefore, it is critical to capture these delayed outcomes in the Stackelberg game formulation.
@@
2 Related Work
2.1 Strategic Manipulation
Though our work primarily lies in proposing a new model for improvement behaviors, the problem settings are also closely related to strategic classification problems (Hardt et al. 2016a; Ben-Porat and Tennenholtz 2017; Dong et al. 2018; Braverman and Garg 2020; Sundaram et al. 2021; Jagadeesan, Mendler-Dünner, and Hardt 2021; Ahmadi et al. 2021; Eilat et al. 2022; Horowitz and Rosenfeld 2023; Miehling et al. 2019). Hardt et al. (2016a) formulated classification problems with strategic manipulation as a Stackelberg game with deterministic cost functions, where the decision maker optimizes classification accuracy based on individuals’ best responses. Afterwards, more sophisticated analytical frameworks were proposed (Dong et al. 2018; Braverman and Garg 2020; Jagadeesan, Mendler-Dünner, and Hardt 2021). Dong et al. (2018) proposed an online algorithm for strategic classification. Braverman and Garg (2020) added randomness to strategic classifiers, while Shao, Blum, and Montasser (2024); Lechner, Urner, and Ben-David (2023) provided a complete analysis of the regret bound for online strategic classification. On the other hand, Sundaram et al. (2021) analyzes the statistical learnability of strategic classification with an SVC classifier. Jagadeesan, Mendler-Dünner, and Hardt (2021) relaxed the standard microfoundations assumption where individuals are perfectly rational to alternative microfoundations where a proportion of individuals may not be strategic, and proposed a noisy response model to tackle the new problem. Zhang et al. (2022) studied the setting where the decision maker and individuals only have knowledge of the feature distributions as random variables. Thus, the strategic manipulation corresponds to a distribution shift and its cost is also a random variable. Eilat et al. (2022) considered the setting where individual responses are dependent and the classifier is learned through graph neural networks. Xie and Zhang (2024a) studied the long-term impacts on welfare and fairness under a sequential strategic learning setting.
2.2 Improvement
However, there are other literature considering improvement behavior (Liu et al. 2019; Rosenfeld et al. 2020; Shavit, Edelman, and Axelrod 2020; Alon et al. 2020; Zhang et al. 2020; Chen, Wang, and Liu 2020; Kleinberg and Raghavan 2020; Bechavod et al. 2021; Ahmadi et al. 2022a, b; Raab and Liu 2021; Xie et al. 2024; Xie and Zhang 2024b; Zhang, Khalili, and Liu 2020). Unlike strategic manipulation, improvement will incur a label change. Liu et al. (2019) studied the conditions where fairness interventions can promote improvement among individuals. Zhang et al. (2020) formulated the label change as a transition matrix where the transition probabilities are deterministic and difficult to estimate. Other works consider both behaviors at the same time. Xie et al. (2024) studied randomness when the agents can both improve and manipulate, while Xie and Zhang (2024b) relaxed the linear assumption on the decision policy and studied the welfare under strategic learning settings. Kleinberg and Raghavan (2020) proposed a mechanism to incentivize individuals to invest in specific features where the individuals have a budget to invest strategically on all features including undesired ones. Their work inherited the classical settings of the Principal-agent model in economics but designed an incentivizing mechanism under a linear machine learning classifier. They modeled manipulation and improvement similarly (linear in efforts) and did not consider the persistent and delayed effects of improvement. By contrast, we first develop a fundamentally different dynamic model to characterize persistent and delayed improvements. Based on the new model, we construct a Stackelberg game to model the interplay between agents and the decision-maker.
Another line of literature focuses on the underlying causal models of strategic classification. Shavit, Edelman, and Axelrod (2020) and Alon et al. (2020) introduced causal inference frameworks into strategic behaviors including manipulation and improvement. Chen, Wang, and Liu (2020) divided the features into immutable features, improvable features and manipulable features and explored linear classifiers which can prevent manipulation and encourage improvement. Jin et al. (2022) also focused on incentivizing improvement and proposed a subsidy mechanism to induce improvement actions and improve social well-being metrics. Barsotti, Kocer, and Santos (2022) conducted several experiments where both improvement and manipulation are possible and both actions incur linear deterministic costs.
2.3 Other Agent Dynamics
In addition to manipulation and improvement, some studies consider different individual behaviors and models. For example, Zhang et al. (2019); Dean et al. (2024a); Hashimoto et al. (2018) focused on participation dynamics where agents at each time decide their participation in decision systems. Perdomo et al. (2020); Hardt, Jagadeesan, and Mendler-Dünner (2022); Jin et al. (2024a) studied performative prediction which is an abstract framework for optimization when the deployed model influences agent population. Yin et al. (2024) studied the setting where the population distribution has an unknown dynamic and reinforcement learning is used to improve long-term fairness. We refer interested readers to survey papers (Zhang and Liu 2021; Dean et al. 2024b) for more examples. Also, our work is related to preference shifts and opinion dynamics in recommendation systems, which we refer to Castellano, Fortunato, and Loreto (2009) as a comprehensive survey. Among the rich set of works, Dean and Morgenstern (2022); Gaitonde, Kleinberg, and Tardos (2021) proposed geometric models for opinion polarization and motivate our work.
@@ 3 Problem Formulation
Decision-maker’s policy.
For an agent with qualification profile $q_{t}$, the decision-maker assesses whether the agent’s profile lines up with the desired qualifications $d$, and makes decision $D_{t}$ based on their similarity $x_{t}:=q_{t}^{T}d$ using a fixed threshold policy $\pi(x_{t})=\textbf{1}(x_{t}\geq\theta)$, i.e., only agents that are sufficiently fit can get accepted. How to choose threshold $\theta$ is discussed in Sec. 5. We assume only agents with initial similarity $x_{0}\geq 0$ are interested in positions and only focus on these candidates.
Although the decision policy introduced above focuses on the similarity between $q_{t}$ and $d$ where qualifications $q_{t}$ are normalized with the same magnitude for all agents, it can be easily extended to settings where the magnitude/strength of skills also matters and may differ across agents. For example, if the decision-maker prefers students with balanced math and English skills and there are two "balanced" students, the decision-maker is likely to prefer the one with higher scores. Therefore, we propose a pre-normalization procedure to take magnitude into account. The idea is to first add an additional dimension to initial qualification profile $q_{0}$, which represents the agent’s unobservable “irrelevant attribute" (all other skills an agent has that are not important for the decision). Thus, we expand the original $q_{0}$ to obtain a $m+1$ dimensional complete qualification profile. Meanwhile, we add a dimension to the ideal qualification profile $d$ with 00 as its value; the new ideal profile becomes $[d;0]$. Then we can make the following natural assumption:
Assumption 3.1.
Assumption 3.1 has been supported by literature in machine learning (Liu, Garg, and Borgs 2022) and social science (Holmstrom and Milgrom 1991). The “irrelevant" dimension demonstrates all other skills that belong to an agent but are not important to the decision. Therefore, competency in relevant/measurable attributes implies weakness in irrelevant/immeasurable attributes and the length of the complete qualification profile stays the same for all agents. With Assumption 3.1 and the distribution of $q_{0}$ as $Q$, we formalize the pre-normalization procedure in Algorithm 1.
Algorithm 1 Pre-normalization procedure
1:  $d=[d,0]$.
3:  for $i\in\{1,\ldots n\}$ do
5:     $q_{0}^{i}=\frac{[q_{0}^{i};z^{i}]^{T}}{K}\in[0,1]^{m+1}$.
6:  end for
Agent qualification dynamics.
We assume agents have information about the ideal profile $d$ (e.g., from application guides, mock interviews). In the beginning, agents with $q_{0}$ can choose to improve their profiles by investing an effort $k\in[0,1]$ to acquire the relevant knowledge, but the effort will have delayed and persistent effects over subsequent time stages. The specific value of $k$ depends on the agent’s utility and will be introduced at the end of this section. Upon investing $k$, the agent’s qualifications $q_{t}$ gradually improve over time based on the following:
@@
Dynamics in (1) model the delayed and persistent impacts of improvement action (i.e., effort $k$). In many real applications, humans acquire knowledge and benefit from repeated practices. They make progress toward the goal gradually, and it takes time to receive the desired outcome from the investment. Indeed, (1) is inspired by the dynamics in Dean and Morgenstern (2022), which was used for modeling preference shifts (details are in Sec. 2.3), where individuals update their opinions/preferences based on their correlations with some influencer (e.g., a political figure) and control the power of the intervention. We believe this is similar to improvement especially when agents improve themselves by imitating some "role models". For example, in a job application scenario, the “influencer” is a current worker who holds information session and introduces her profile $d$. Then the agents strive to mimic the profile of $d$ by updating $q_{t}$. The imitating nature of improvement is well justified in many works (e.g., Raab and Liu (2021); Zhang et al. (2022)), making agent improvement suitable to be modeled in a similar fashion to concept drift/preference shift. Thus, we use (1) to model the evolution of agents’ (pre-normalized) qualifications. Based on Prop. 1 of Dean and Morgenstern (2022), we know that $q_{t}$ converges under dynamics, as formally stated in Lemma 3.2 below.
@@
Lemma 3.2 suggests that any agent eventually becomes an ideal candidate with a perfectly aligned profile (i.e., $x_{t}=q_{t}^{T}d=1$), as long as he/she is interested in the position ($x_{0}\geq 0$) and willing to make an effort ($k>0$). The only difference among agents is the speed of convergence: it takes less time for agents who are more qualified at the beginning (i.e., larger $x_{0}$) and/or make more effort (i.e., larger $k$) to become ideal and get accepted. Note that our work focuses on agent’s improvement behavior with persistent and delayed effects. Although the model is presented in a simplified setting where only a one-step effort is made by the agents at the beginning, it can capture more complicated scenarios where agents repeatedly exert efforts multiple times until they reach the target. Each effort has persistent effects on improving the qualification. Specifically, suppose each agent at time $t$ can exert an effort $k_{t}$ and the agent’s qualification improves based on $\widetilde{q}_{t+1}=q_{t}+\sum_{\tau=0}^{t}k_{\tau}\cdot q_{t}^{T}d\cdot d$ with $q_{t+1}=\frac{\widetilde{q}_{t+1}}{\|\widetilde{q}_{t+1}\|_{2}}$ (i.e., every time the agent improves from all accumulated efforts $\sum_{\tau=0}^{t}k_{\tau}\in[0,1]$ he/she invested so far). For this new dynamics, the overall impacts of these efforts $\{k_{t}\}_{t\geq 0}$ on improving agent qualification can be equivalently characterized by the dynamics (1) with some one-step effort. That is, there exists an effort $k^{*}\in[0,1]$ such that investing $k^{*}$ once at the beginning has the same impact on $\lim_{t\to\infty}q_{t}$ as investing a sequence of efforts $\{k_{t}\}_{t\geq 0}$ over time. We provide more detailed discussion on this in App. A. We also discuss a special case where the effect of $k$ is decreasing over time and provide further convergence analysis (Thm. A.1) in App. A.
Agent’s utility & action.
Because it takes time for agents to receive rewards (i.e., get accepted) for their efforts, they may not have incentives to invest if there is a long delay. In practice, people may be more attracted to investments with immediate rewards than delayed rewards, or they may simply not have enough time to wait. For example, students only have limited time to prepare for college applications; credit card applicants may not have incentives to improve their credit scores and wait to get approval for a specific credit card when there are many instant-approval cards on the market.
@@
That is, the utility is the exponentially discounted reward an agent receives from the acceptance minus the effort. $r>0$ is the discounting factor. Note that the discounted utility model [^1] Under exponential discounting function, the agent’s reward diminishes at a constant rate (Grüne-Yanoff 2015). Our model can also adopt other discounting functions (e.g., hyperbolic discounting) for settings when the agent’s reward decreases inconsistently. The qualitative results of this paper still remain the same. has been widely used in literature such as reinforcement learning (Kaelbling, Littman, and Moore 1996), finance (Meier and Sprenger 2013), and economics (Krahn and Gafni 1993; Samuelson 1937).
@@
Stackelberg game.
We model the strategic interplay between the decision-maker and agents as a Stackelberg game, which consists of two stages: (i) the decision-maker first publishes the optimal acceptance threshold $\theta$ (details are in Sec. 5); (ii) agents after observing the threshold take actions to maximize their utilities as given in (5).
Manipulation & forgetting.
The model formulated above has two implicit assumptions: (i) agents are honest and they improve their qualifications by making actual efforts; (ii) once agents make a one-time effort $k$ to acquire the knowledge, they never forget and can repeatedly leverage this knowledge to improve their profiles based on (1). However, these assumptions may not hold. In practice, agents may fool the decision-maker by directly manipulating $x_{t}$ to get accepted without improving actual $q_{t}$, e.g., people cheat on exams or interviews to get accepted. Moreover, the knowledge agents acquired at the beginning may not be sufficient to ensure repeated improvements. To capture these, we further extend the above model to two settings:
@@
Objective.
In this paper, we study the above interactions between decision-maker and agents. We aim to understand (i) under what conditions agents have incentives to improve their qualifications; (ii) how to design the optimal policy to incentivize the largest improvements inside the agent population; (iii) how the agents would behave when they have both options of manipulation and improvement, and under what conditions agents prefer improvement over manipulation; (iv) how the forgetting mechanism affects agent’s behavior and long-term qualifications.